# Twin Deformable Point Convolutions for Point Cloud Semantic Segmentation in Remote Sensing Scenes

### v1

@@ IV-B2 Qualitative Analysis
ISPRS Vaihingen 3D: From Fig. 5, we provide qualitative visualization results of semantic segmentation of TDConvs on the ISPRS Vaihingen 3D test set. Through observation and analysis, our proposed TDConvs achieves superior results in overall segmentation results. Specifically, it can be found from the figure that the proposed TDConvs can perform better and accurate classification regardless of large-scale areas or small areas. In addition, in the locally enlarged area, especially small objects and objects with obvious height differences can be well classified. In addition, our TDConvs also has superior segmentation capabilities for sparse objects such as powerlines.
@@
LASDU: From Fig. 6, we give the qualitative analysis results of semantic segmentation of TDConvs on the LASDU dataset. In comparison with Ground Truth (GT), the prediction results of TDConvs are basically consistent with GT and achieve superior segmentation results on objects such as roads and buildings. By changing the visual angle of part of the area, it can be found that buildings and trees with obvious height differences are well distinguished, which is challenging in the segmentation task of large-scale point cloud scenes.
TABLE III: Ablation experimental analysis of various components of the proposed TDConvs.
@@
(a) Performance of baseline method. (b) Performance of baseline which adds CyDConv without MFL. (c) Performance of baseline which adds CyDConv with MFL. (d) Performance of baseline with both CyDConv and SpDConv. ‘$\Delta$’ represents the increment of performance.
CyDConv w/o MFL CyDConv w/ MFL SpDConv OA mF1 $\Delta$(mF1) $\sum\Delta$
(a) 82.2 68.6 - -
(b) 83.6 70.8 +2.2 2.2
(c) 84.0 72.2 +1.4 3.6
(d) 84.5 73.4 +1.2 4.8
TABLE IV: Verification experiments on the grid scale of CyDConv and SpDConv.
@@
The first four columns represent the resolution scales of maps at different levels, and the fifth column represents the resolution scale of the volume (remaining unchanged).
@@
In addition, we also compared the prediction results with other methods, including the popular PointConv[7] and RFFS-Net[1]. Through comparison, it can be found that the TDConvs we proposed has achieved superior segmentation results in road segmentation, objects with large height differences such as buildings and trees, and highly similar objects such as the ground and low vegetation. However, PointConv[7] and RFFS-Net[1] have poor performance in these two fields. In this case, there is no advantage for these two methods. The reason why our TDConvs achieve such superior results is that the proposed TDConvs can well model the latitude, longitude, and altitude characteristics of point clouds in large-scale scenes, which is greatly beneficial to the segmentation of different instance objects, especially for larger instances that are difficult to separate and with large height differences or high similarity.
IV-C Ablation Studies
IV-C1 The Effect of CyDConv
CyDConv is proposed to explicitly model the arrangement characteristics of point clouds in longitude and latitude in large-scale scenes. As shown in Table III, based on (a) baseline, we added CyDConv but did not introduce MFL operation, and the network performance improved from 82.2% OA and 68.6% mF1 to 83.6% OA and 70.8% mF1. mF1 increased by 2.2 percent over the baseline model, which is very advantageous in the semantic segmentation task of large-scale point clouds. This is because the CyDConv we proposed can provide explicit longitude and latitude modeling for point cloud data obtained from the aerial perspective, which helps to improve segmentation accuracy.
IV-C2 The Effect of MFL
Multi-scale neighbor point feature extraction can simultaneously extract features of point sets in different neighbor areas of the center point, promoting the network’s fine feature learning of objects of different scales. The introduction of MFL can realize this function well. As shown in Table III, based on (b), we added MFL operation to CyDConv, which improved the network performance from the original 83.6% OA and 70.8% mF1 to the current 84.0% OA and 72.2% mF1, respectively. OA and mF1 increased by 0.4% and 1.4%. With MFL, the network can extract features from various instances at different scales, promoting the improvement of the overall network segmentation performance. The reason is that the selection of different neighboring areas allows the center point to have different receptive fields, which is helpful for feature learning of objects of multiple scales in large-scale scenes.
TABLE V: Verification experiments on the grid scale of CyDConv and SpDConv. The first four columns represent the resolution scales of maps at different levels (remaining unchanged), and the fifth column represents the resolution scale of the volume.
@@ IV-C2 The Effect of MFL
IV-C3 The Effect of SpDConv
IV-D Deep Analysis
IV-D1 Grid Scale of CyDConv and SpDConv
Since the proposed CyDConv and SpDConv are both based on map generation and volume generation in small resolution, the selection of the resolution is a significant process that determines network performance. From Table IV, we explore the impact of CyDConv’s different resolution maps on overall performance by changing the resolution scales of maps at different levels. Under the premise that the volume resolution remains unchanged, different map resolutions have different effects on the segmentation performance of our TDConvs. In particular, as the resolution of each layer is continuously increased, the segmentation performance shows a trend of first increasing and then decreasing. Excessive resolution does not bring better performance, but increases the amount of calculation and parameters. Therefore, TDConvs finally chose $[40\times 40,20\times 20,10\times 10,5\times 5]$. In addition, from Table V, while ensuring that the resolution of the map remains unchanged, changing the resolution of the volume will also have different effects on the segmentation performance of our TDConvs. The increase in volume resolution causes a downward trend in performance, so we finally choose $[40\times 40\times 5]$ as the volume resolution.
IV-D2 Sampling Number of CyDConv and SpDConv
TABLE VI: Analysis of the number of offset points in 2D map and 3D volume. ‘offsets in 2D map’ and ‘offsets in 3D volume’ represent the number of offsets in the 2D map space and the number of offsets in the 3D volume space, respectively.
offsets in 2D Map offsets in 3D Volume OA mF1
2 4 81.6 71.6
4 8 84.5 73.4
8 16 84.3 72.2
16 32 82.2 71.3
Within map and volume, different numbers of offset points have a certain impact on network performance. We set different numbers of sampling points of CyDConv and SpDConv in the experiments to determine the parameters. As shown in Table VI, the network can achieve optimal segmentation performance by setting the number of offsets in 2D map and the number of offsets in 3D volume to 4 and 8, respectively. The combination of 8 and 16 also achieved good performance, but it brought more calculations and the performance was lower than the combination of 4 and 8. An appropriate number of sampling points can promote CyDConv and SpDConv to better learn the longitude, latitude, and altitude information of surrounding points, which will help improve the overall network segmentation performance.
IV-D3 Number of Nearest Neighbor Sampling Points in MFL
TABLE VII: Combination analysis of the number of nearest neighbor points in the MFL module. $\rm T_{1}$, $\rm T_{2}$, and $\rm T_{3}$ represent the number of points in the three neighboring areas, respectively.
@@ IV-D3 Number of Nearest Neighbor Sampling Points in MFL
The construction of neighboring points at different levels can promote the improvement of the network’s perception of different areas. The introduction of MFL enables the network to learn neighbor features of different radii within the point set during the feature extraction process. To better determine the number of aggregation points for different neighbor points, we conducted ablation experiments on the number of different aggregation points. As shown in Table VII, we determine the final MFL neighbor point combination by setting different $\rm T_{1}$, $\rm T_{2}$, and $\rm T_{3}$ values. It is obvious that in the combination $[16,32,64]$, the network can make full use of points in different sensing areas to perform feature aggregation at different levels. Although the network combined as $[32,64,128]$ also achieved good performance, more sampling points brought unnecessary calculations.
Figure 8: Parametric analysis of the proposed TDConvs. The abscissa is the parameter amount, and the ordinate is the mF1 performance of segmentation. The closer the point is to the upper left corner, the more efficient the network is.
IV-D4 Paramaters Analysis
@@
The parameter amount and performance of the network are important indicators to evaluate the efficiency of the model. In practical applications, low parameters and high segmentation performance are necessary. As shown in Fig. 8, we compared the number of parameters and segmentation performance of existing advanced methods and the proposed TDConvs and gave the relevant two-dimensional diagrams. The abscissa of the figure is the parameters of the models, and the ordinate is the mF1 indicator of segmentation. Obviously, the smaller the abscissa and the larger the ordinate represent the higher the efficiency of the model and the better the performance. It can be concluded from the figure that the TDConvs method we proposed is located in the upper left corner of the figure, which shows that we have achieved the best segmentation performance while ensuring the low number of parameters of our TDConvs, which effectively guarantees the practical application of lightweight.
IV-D5 Statistical Analysis
@@
Analysis of statistical characteristics of network performance is necessary. As shown in Fig. 9, we counted the category-by-category performance distribution of the baseline method and the TDConvs proposed in this paper. Through analysis, we found that the segmentation performance of our TDConvs is better than the baseline in all categories, and its performance is particularly outstanding in individual categories, such as the fence category and the tree category. From the statistical analysis, we can conclude that in point cloud segmentation of large-scale scenes, it is necessary to explicitly model its longitude, latitude, and altitude information, which can greatly promote the overall segmentation performance.
@@
Figure 9: Category statistical analysis of the proposed TDConvs and baseline. The abscissa is each category, and the ordinate is the performance corresponding to each category.
V Conclusion
@@
We propose Twin Deformable point Convolutions (TDConvs) to explicitly model the geographic information characteristics of longitude, latitude, and altitude of large-scale aerial point clouds. First, to learn longitude and latitude geographic information features, we proposed Cylinder-wise Deformable point Convolution (CyDConv), which performs adaptive aggregation of features by learning deformable sampling points of point clouds in a two-dimensional map space to improve the learning ability of longitude and latitude features. In addition, altitude learning is significant in large-scale point cloud processing in the remote sensing field. To this end, we introduce Sphere-wise Deformable point Convolution (SpDConv), which promotes the network for the fine-grained classification of highly different and highly similar instances through adaptively learning the altitude features of points in the three-dimensional volume space. By introducing TDConvs, the network can efficiently learn geographical information features such as the latitude, longitude, and altitude information of point clouds, which helps in fine classification in actual large-scale scenarios and promotes implementation in practical applications. TDConvs provide a new perspective for the semantic segmentation task of large-scale point clouds in the remote sensing field.

### v2

@@ 4.2.2 Qualitative Analysis
ISPRS Vaihingen 3D: From Fig. 4.2.1, we provide qualitative visualization results of semantic segmentation of TDConvs on the ISPRS Vaihingen 3D test set. Through observation and analysis, our proposed TDConvs achieves superior results in overall segmentation results. Specifically, it can be found from the figure that the proposed TDConvs can perform better and accurate classification regardless of large-scale areas or small areas. In addition, in the locally enlarged area, especially small objects and objects with obvious height differences can be well classified. In addition, our TDConvs also has superior segmentation capabilities for sparse objects such as powerlines.
Fig. 6.: Segmentation results of the proposed TDConvs on the LASDU test set. The upper part is the ground truth of the dataset, and the lower part is the segmentation results of the proposed TDConvs.
Fig. 7.: Comparison of the segmentation results of the proposed TDConvs with other methods on the LASDU test set. The first column is the segmentation results of PointConv[53], the second column is the segmentation results of RFFS-Net[32], and the third column is the segmentation results of the proposed TDConvs.
Fig. 8.: Comparison of the segmentation results of the proposed TDConvs with other methods on the DFC2019 test set. The first column is the Ground Truth, the second column is the segmentation results of PointConv[53], the third column is the segmentation results of RFFS-Net[32], the fourth column is the segmentation results of VD-LAB[23], and the fifth column is the segmentation results of the proposed TDConvs.
@@
LASDU: From Fig. 4.2.2, we give the qualitative analysis results of semantic segmentation of TDConvs on the LASDU dataset. In comparison with Ground Truth (GT), the prediction results of TDConvs are basically consistent with GT and achieve superior segmentation results on objects such as roads and buildings. By changing the visual angle of part of the area, it can be found that buildings and trees with obvious height differences are well distinguished, which is challenging in the segmentation task of large-scale point cloud scenes.
@@
Table 4: Ablation experimental analysis of various components of the proposed TDConvs. (a) Performance of baseline method. (b) Performance of baseline which adds CyDConv without MFL. (c) Performance of baseline which adds CyDConv with MFL. (d) Performance of baseline with both CyDConv and SpDConv. ‘$\Delta$’ represents the increment of performance or parameters.
CyDConv w/o MFL CyDConv w/ MFL SpDConv OA mF1 $\Delta$(mF1) $\sum\Delta$ Params. $\Delta$(Params.)
(a) 82.2 68.6 - - 1.42M -
(b) 83.6 70.8 +2.2 2.2 2.24M +0.82M
(c) 84.0 72.2 +1.4 3.6 3.28M +1.04M
(d) 84.5 73.4 +1.2 4.8 3.31M +0.03M
In addition, we also compared the prediction results with other methods in Fig. 4.2.2, including the popular PointConv[53] and RFFS-Net[32]. Through comparison, it can be found that the TDConvs we proposed has achieved superior segmentation results in road segmentation, objects with large height differences such as buildings and trees, and highly similar objects such as the ground and low vegetation. However, PointConv[53] and RFFS-Net[32] have poor performance in these two fields. In this case, there is no advantage for these two methods. The reason why our TDConvs achieve such superior results is that the proposed TDConvs can well model the latitude, longitude, and altitude characteristics of point clouds in large-scale scenes, which is greatly beneficial to the segmentation of different instance objects, especially for larger instances that are difficult to separate and with large height differences or high similarity.
DFC2019: Fig. 4.2.2 shows the visual comparisons of existing methods (including PointConv, RFFS-Net, and VD-LAB) and our method on the DFC2019 dataset. As can be seen from the figure, the method in this paper has superior segmentation performance compared to other methods, which is clearly shown in the blue boxes. In particular, in different areas, including building areas and natural feature areas, our TDConvs can achieve superior segmentation in details and is consistent with the Ground Truth. The reason is that the explicit modeling of longitude, latitude, and altitude in this paper can have good feature extraction capabilities for such large-scale point clouds to obtain the best segmentation effect.
4.3 Ablation Studies
4.3.1 The Effect of CyDConv
CyDConv is proposed to explicitly model the arrangement characteristics of point clouds in longitude and latitude in large-scale scenes. As shown in Table 4, based on (a) baseline, we added CyDConv but did not introduce MFL operation, and the network performance improved from 82.2% OA and 68.6% mF1 to 83.6% OA and 70.8% mF1. mF1 increased by 2.2 percent over the baseline model, which is very advantageous in the semantic segmentation task of large-scale point clouds. This is because the CyDConv we proposed can provide explicit longitude and latitude modeling for point cloud data obtained from the aerial perspective, which helps to improve segmentation accuracy.
4.3.2 The Effect of MFL
Multi-scale neighbor point feature extraction can simultaneously extract features of point sets in different neighbor areas of the center point, promoting the network’s fine feature learning of objects of different scales. The introduction of MFL can realize this function well. As shown in Table 4, based on (b), we added MFL operation to CyDConv, which improved the network performance from the original 83.6% OA and 70.8% mF1 to the current 84.0% OA and 72.2% mF1, respectively. OA and mF1 increased by 0.4% and 1.4%. With MFL, the network can extract features from various instances at different scales, promoting the improvement of the overall network segmentation performance. The reason is that the selection of different neighboring areas allows the center point to have different receptive fields, which is helpful for feature learning of objects of multiple scales in large-scale scenes.
4.3.3 The Effect of SpDConv
@@
Table 5: Verification experiments on the grid scale of CyDConv and SpDConv. The first four columns represent the resolution scales of maps at different levels, and the fifth column represents the resolution scale of the volume (remaining unchanged).
@@
4.4 Deep Analysis
4.4.1 Grid Scale of CyDConv and SpDConv
Since the proposed CyDConv and SpDConv are both based on map generation and volume generation in small resolution, the selection of the resolution is a significant process that determines network performance. From Table 5, we explore the impact of CyDConv’s different resolution maps on overall performance by changing the resolution scales of maps at different levels. Under the premise that the volume resolution remains unchanged, different map resolutions have different effects on the segmentation performance of our TDConvs. In particular, as the resolution of each layer is continuously increased, the segmentation performance shows a trend of first increasing and then decreasing. Excessive resolution does not bring better performance, but increases the amount of calculation and parameters. Therefore, TDConvs finally chose $[40\times 40,20\times 20,10\times 10,5\times 5]$. In addition, from Table 7, while ensuring that the resolution of the map remains unchanged, changing the resolution of the volume will also have different effects on the segmentation performance of our TDConvs. The increase in volume resolution causes a downward trend in performance, so we finally choose $[40\times 40\times 5]$ as the volume resolution.
4.4.2 Sampling Number of CyDConv and SpDConv
Table 6: Analysis of the number of offset points in 2D map and 3D volume. ‘offsets in 2D map’ and ‘offsets in 3D volume’ represent the number of offsets in the 2D map space and the number of offsets in the 3D volume space, respectively.
offsets in 2D Map offsets in 3D Volume OA mF1
2 4 81.6 71.6
4 8 84.5 73.4
8 16 84.3 72.2
16 32 82.2 71.3
Within map and volume, different numbers of offset points have a certain impact on network performance. We set different numbers of sampling points of CyDConv and SpDConv in the experiments to determine the parameters. As shown in Table 6, the network can achieve optimal segmentation performance by setting the number of offsets in 2D map and the number of offsets in 3D volume to 4 and 8, respectively. The combination of 8 and 16 also achieved good performance, but it brought more calculations and the performance was lower than the combination of 4 and 8. An appropriate number of sampling points can promote CyDConv and SpDConv to better learn the longitude, latitude, and altitude information of surrounding points, which will help improve the overall network segmentation performance.
Table 7: Verification experiments on the grid scale of CyDConv and SpDConv. The first four columns represent the resolution scales of maps at different levels (remaining unchanged), and the fifth column represents the resolution scale of the volume.
@@ 4.4.2 Sampling Number of CyDConv and SpDConv
4.4.3 Number of Nearest Neighbor Sampling Points in MFL
Table 8: Combination analysis of the number of nearest neighbor points in the MFL module. $\rm T_{1}$, $\rm T_{2}$, and $\rm T_{3}$ represent the number of points in the three neighboring areas, respectively.
@@ 4.4.3 Number of Nearest Neighbor Sampling Points in MFL
The construction of neighboring points at different levels can promote the improvement of the network’s perception of different areas. The introduction of MFL enables the network to learn neighbor features of different radii within the point set during the feature extraction process. To better determine the number of aggregation points for different neighbor points, we conducted ablation experiments on the number of different aggregation points. As shown in Table 8, we determine the final MFL neighbor point combination by setting different $\rm T_{1}$, $\rm T_{2}$, and $\rm T_{3}$ values. It is obvious that in the combination $[16,32,64]$, the network can make full use of points in different sensing areas to perform feature aggregation at different levels. Although the network combined as $[32,64,128]$ also achieved good performance, more sampling points brought unnecessary calculations.
4.4.4 Paramaters Analysis
@@
The parameter amount and performance of the network are important indicators to evaluate the efficiency of the model. In practical applications, low parameters and high segmentation performance are necessary. As shown in Fig. 4.4.4, we compared the number of parameters and segmentation performance of existing advanced methods and the proposed TDConvs and gave the relevant two-dimensional diagrams. The abscissa of the figure is the parameters of the models, and the ordinate is the mF1 indicator of segmentation. Obviously, the smaller the abscissa and the larger the ordinate represent the higher the efficiency of the model and the better the performance. It can be concluded from the figure that the TDConvs method we proposed is located in the upper left corner of the figure, which shows that we have achieved the best segmentation performance while ensuring the low number of parameters of our TDConvs, which effectively guarantees the practical application of lightweight.
Fig. 9.: Parametric analysis of the proposed TDConvs. The abscissa is the parameter amount, and the ordinate is the mF1 performance of segmentation. The closer the point is to the upper left corner, the more efficient the network is.
As shown in Table 4, we give the calculation results of the number of parameters for (a) baseline, (b) baseline with CyDConv, (c) baseline with CyDConv and MFL, and (d) baseline with CyDConv, MFL, and SpDConv. From the given data, it can be seen that our TDConvs has a very large advantage in the overall number of parameters, and can achieve high segmentation performance while ensuring a low number of parameters. For each module, the increase in the number of parameters is within an acceptable range and is negligible. For CyDConv, method (b) increases the number of parameters by 0.82M compared to method (a). Furthermore, after adding MFL, the number of parameters increases from 2.24M in method (b) to 3.28M in method (c), an increase of 1.04M. Finally, SpDConv increases the number of parameters by 0.03M. These are tiny increases in the number of parameters and are within the acceptable range for actual network deployment. It can be seen that the method proposed in this paper ensures high segmentation performance while ensuring low parameter increase.
4.4.5 Statistical Analysis
@@
Analysis of statistical characteristics of network performance is necessary. As shown in Fig. 4.4.5, we counted the category-by-category performance distribution of the baseline method and the TDConvs proposed in this paper. Through analysis, we found that the segmentation performance of our TDConvs is better than the baseline in all categories, and its performance is particularly outstanding in individual categories, such as the fence category and the tree category. From the statistical analysis, we can conclude that in point cloud segmentation of large-scale scenes, it is necessary to explicitly model its longitude, latitude, and altitude information, which can greatly promote the overall segmentation performance.
@@
Fig. 10.: Category statistical analysis of the proposed TDConvs and baseline. The abscissa is each category, and the ordinate is the performance corresponding to each category.
5 Conclusion
@@
We proposed Twin Deformable point Convolutions (TDConvs) to explicitly model the geographic information characteristics of longitude, latitude, and altitude of large-scale aerial point clouds. First, we proposed Cylinder-wise Deformable point Convolution (CyDConv), which performed adaptive aggregation of features by learning deformable sampling points of point clouds in a two-dimensional map space to improve the learning ability of longitude and latitude features. In addition, we introduced Sphere-wise Deformable point Convolution (SpDConv), which promoted the network for the fine-grained classification of highly different and highly similar instances through adaptively learning the altitude features of points in the three-dimensional volume space. By introducing TDConvs, the network efficiently learned geographical information features such as the latitude, longitude, and altitude information of point clouds, which helped in fine classification in actual large-scale scenarios and promoted implementation in practical applications. TDConvs provided a new perspective for the semantic segmentation task of large-scale point clouds in the remote sensing field.
Future Work. In future research, we will focus on two aspects: multi-source data fusion and foundation models of point cloud. First, the data of various existing sensors are developing rapidly. Single-modality data often has certain limitations and unreliable results for segmentation tasks, while the fusion of multi-source data can solve the unreliability of prediction results and improve the robustness of interpretation. Therefore, how to fuse point cloud data with other types of remote sensing data (such as remote sensing images) to improve the accuracy of data interpretation is the focus of our future research. In addition, the complexity and diversity of point cloud data require efficient and flexible processing methods. Traditional methods are usually effective for specific types of data or specific tasks, but difficult to extend to other scenarios. The large foundation model can learn general feature representations on large-scale and diverse data sets through pre-training, so it has a wider applicability. Therefore, how to build a point cloud foundation model under large-scale data is an urgent need for research in the field of remote sensing.
6 Acknowledgments